OpenStudy (anonymous):

in the isosceles triangle ABC, AB is congruent to CB. if AB=2x+17, CB=4x-13,and AC=2x+16, find the value of x.

OpenStudy (arabpride):

AB is congruent to CB AB = CB 2x + 17 = 4x - 13 -2x -2x 17 = 2x - 13 +13 +13 30 = 2x /2 /2 15 = x or x = 15

OpenStudy (arabpride):

Does that make sense??

OpenStudy (anonymous):

yes! thank you!

OpenStudy (arabpride):

You're Very Welcome *^.^* If you have another one question im happy to help!!